OpenStudy (enigma123):

Session 77 lecture question: How would you evaluate the triple integral in spherical coordinates in his example? (how to integrate sec^3(theta)?)

OpenStudy (phi):

can you post the integral?

OpenStudy (enigma123):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi/4}\int\limits_{1/\sqrt{2}\sec{\phi}}^{1} {\rho}^2\sin{\phi} d{\rho}d{\phi}d{\theta}\] It's the volume of the portion of the unit sphere above z = 1/sqrt(2)

OpenStudy (phi):

the inner gives \[ \frac{\rho ^3}{3} \bigg|_{\sqrt{2}\sec \phi}^1= \frac{1}{3} - \frac{1}{6\sqrt{2}} \cos^{-3} \phi\] the middle is \[ \int_0^\frac{\pi}{4} \frac{1}{3} \sin \phi - \frac{1}{6\sqrt{2}}\cos^{-3} \sin \phi \ d\phi\] the first term gives \[ -\frac{1}{3} \cos \phi \bigg|_0^\frac{\pi}{4} = - \frac{1}{3\sqrt{2}} +\frac{1}{3}\] the second term is an example of u^(-3) du where u= cos x and du = - sin x dx it integrates to \[ \frac{1}{6\sqrt{2}} \frac{\cos^{-2}\phi}{-2} \bigg|_0^\frac{\pi}{4} \\ = -\frac{1}{12\sqrt{2}}\left( \frac{1}{\frac{1}{2}} - 1\right)= -\frac{1}{12\sqrt{2}} \] the two terms add up to \[ \frac{1}{3} - \frac{1}{3\sqrt{2}}-\frac{1}{12\sqrt{2}}\\ = \frac{1}{3} - \frac{5}{12\sqrt{2}} \] finally, the last integral \[ \int_0^{2\pi} \ d \theta = 2 \pi \] and the answer is \[ \frac{2 \pi}{3} - \frac{5\pi}{6\sqrt{2}}\]

OpenStudy (phi):

oops: I noticed that I show \[ \sqrt{2} \sec \phi\] as the lower bound on the first integral. It shoud read \[ \frac{1}{\sqrt{2}}\sec \phi\] The calculations use this value, so the result is correct.

OpenStudy (csbuja):

Phi, you are awesome.